For the two unrelated situations below, compute the unknown amounts indicated by the letters (a) and (b).

(1)

(2)

Beginning:

Assets

$42,000

$61,000

Liabilities

18,000

24,000

Ending:

Assets

48,000

68,000

Liabilities

15,000

(b)

During the year:

Stock issued

(a)

4,000

Revenues

28,000

45,000

Dividends paid

5,000

3,000

Expenses

16,000

37,000
Answer #1

A) $ 2,000

B) $ 22,000

Working:

A. Ending Equity = Beginning Equity + Stock Issued + Revenue - Expenses -Dividends Paid

[48000-15000] = [42000-18000] + Stock Issued + 28000 - 16000 - 5000

33000 = 31000 + Stock Issued

Stock Issued = $ 2000 [33000-31000]

B. Ending Liabilities = Ending Assets - Ending Equity

Ending Liabilities = 68000 -{ [Beginning Equity + Stock Issued + Revenue - Expenses -Dividends Paid ]

Ending Liabilities = 68000 -{ [61000-24000] + 4000+45000-37000-3000]}

Ending Liabilities = 68000 -46000 = 22000